### v1

# DPMambaIR: All-in-One Image Restoration via Degradation-Aware Prompt State Space Model

Abstract.
All-in-One image restoration aims to address multiple image degradation problems using a single model, significantly reducing training costs and deployment complexity compared to traditional methods that design dedicated models for each degradation type. Existing approaches typically rely on Degradation-specific models or coarse-grained degradation prompts to guide image restoration. However, they lack fine-grained modeling of degradation information and face limitations in balancing multi-task conflicts. To overcome these limitations, we propose DPMambaIR, a novel All-in-One image restoration framework. By integrating a Degradation-Aware Prompt State Space Model (DP-SSM) and a High-Frequency Enhancement Block (HEB), DPMambaIR enables fine-grained modeling of complex degradation information and efficient global integration, while mitigating the loss of high-frequency details caused by task competition. Specifically, the DP-SSM utilizes a pre-trained degradation extractor to capture fine-grained degradation features and dynamically incorporates them into the state space modeling process, enhancing the model’s adaptability to diverse degradation types. Concurrently, the HEB supplements high-frequency information, effectively addressing the loss of critical details, such as edges and textures, in multi-task image restoration scenarios. Extensive experiments on a mixed dataset containing seven degradation types show that DPMambaIR achieves the best performance, with 27.69dB and 0.893 in PSNR and SSIM, respectively. These results highlight the potential and superiority of DPMambaIR as a unified solution for All-in-One image restoration.
All-in-One Image Restoration, State Space Model, Prompt Learning
1. Introduction
Image restoration is a fundamental task in computer vision, aiming to recover high-quality images from degraded inputs. Recent advances in deep learning have significantly improved performance in Degradation-specific restoration problems, such as image denoising (Kim et al., 2024; Zeng et al., 2024; Ryou et al., 2024), deblurring (Chen et al., 2024b; Wu et al., 2024a; Mao et al., 2024), deraining(Chen et al., 2024a, 2023; Ba et al., 2022; Peng et al., 2020, 2021; He et al., 2024b), dehazing(Zhang et al., 2024e; Wang et al., 2024c; Zheng et al., 2023; Wu et al., 2023) and low-light enhancement(Ma et al., 2022; Cai et al., 2023; Hai et al., 2023; Wang et al., 2022, 2023b). However, traditional methods typically rely on Degradation-specific models explicitly tailored to individual degradation types. These approaches often require prior knowledge of the degradation type, which limits their practicality in real-world scenarios, such as autonomous driving and nighttime surveillance, where diverse and unknown degradations frequently coexist.
Figure 1. Comparison with SOTA All-in-One Image Restoration Methods on mixed Dataset. Our DPMambaIR achieves consistently better performance.
To address this limitation, recent studies explore All-in-One image restoration frameworks that unify multiple restoration tasks into a single model. Existing methods can be broadly categorized into Mixture-of-Experts (MoE)-based and Prompt-based approaches. MoE-based methods, such as MoFME (Zhang et al., 2024b), dynamically route tasks to specialized experts using an uncertainty-aware router, enabling effective learning of diverse expert functions. However, these methods suffer from high computational costs, increased inference complexity as the number of experts grows, and lack explicit degradation modeling, limiting adaptability to unseen or complex degradations. Prompt-based methods, like OneRestore (Guo et al., 2024a), guide restoration using degradation prompts, integrating scene descriptors via cross-attention mechanisms for flexible and adaptive strategies. While effective, these methods rely on coarse prompts that fail to capture nuanced degradation characteristics and treat prompts as static inputs, limiting their ability to adapt to varying degradation intensities or specific requirements, leading to suboptimal results.
While these approaches reduce computational cost and deployment complexity by leveraging feature sharing and multi-task learning, two key challenges remain unresolved. First, existing models lack fine-grained degradation awareness and often rely on coarse degradation labels to guide the restoration process. For example, images with varying degrees of haze are typically assigned a single haze label, failing to capture the nuanced variations in degradation levels. This simplification leads to suboptimal restoration performance, as the model cannot adapt to the diverse characteristics of different degradations. Second, current methods struggle to balance the competing requirements of varying restoration tasks. For instance, training low-light enhancement, which focuses on restoring low-frequency contrast information, alongside deblurring, which prioritizes the recovery of fine local details, often results in task competition. This inherent task competition makes it challenging to optimize a single network for all restoration tasks simultaneously and remains a fundamental challenge for All-in-One image restoration frameworks.
To address these challenges, we propose a novel All-in-One image restoration framework, DPMambaIR. The framework is built upon a Degradation-aware Prompt State Space Model (DP-SSM) and a lightweight High-frequency Enhancement Block (HEB). Specifically, DP-SSM employs a pre-trained degradation extractor to extract degradation embeddings, which are used to modulate the state transition, input, and output matrices of the state space equations. This enables fine-grained degradation modeling and efficient global degradation integration. Meanwhile, HEB mitigates the loss of high-frequency details caused by task competition among different restoration tasks. Extensive experiments on a mixed dataset containing seven degradation types demonstrate the superiority of DPMambaIR. As shown in Figure 1, DPMambaIR achieves state-of-the-art performance in terms of PSNR and SSIM, outperforming existing All-in-One methods such as AdaIR and OneRestore. Moreover, DPMambaIR achieves competitive results across individual tasks, including deraining, low-light enhancement, deblurring, and dehazing, showcasing its potential as an effective and robust All-in-One image restoration solution.
The contributions of this paper can be summarized:
  • $\bullet$
    We propose a novel degradation extractor capable of capturing fine-grained degradation features from complex degraded images. These features provide critical priors to improve the performance of All-in-One image restoration.
  • $\bullet$
    We propose a Degradation-aware Prompt State Space Model (DP-SSM) that incorporates fine-grained degradation features into the state space modeling process, enabling both global degradation modeling and fine-grained degradation awareness to effectively handle diverse degradation types within a unified framework.
  • $\bullet$
    Extensive experiments on a mixed dataset containing seven types of degradation demonstrate that DPMambaIR achieves state-of-the-art performance on PSNR and SSIM metrics, outperforming existing methods. The results validate the effectiveness and superiority of the proposed approach.
Figure 2. The overall framework of DPMambaIR consists of a shallow feature extractor and multiple DPSS blocks, with an upsampler implemented by pixel shuffle and a downsampler by pixel unshuffle. Additionally, a degradation extractor is included to extract degradation embeddings, which are then inserted into the DPSS blocks.
2. Related Work
2.1. All-in-One Image Restoration
Image restoration is a fundamental task in computer vision, aiming to recover high-quality images from degraded inputs. Traditional methods, such as Dark Channel Prior (He et al., 2010) and Color Line Prior (Fattal, 2014), rely on handcrafted priors to model degradation but struggle with complex scenarios due to predefined assumptions. The rise of deep learning (LeCun et al., 2015) has significantly advanced tasks like such as image super-resolution(Peng et al., 2024c, d, 2025c; Wang et al., 2024b; Saharia et al., 2022; Gao et al., 2023), image denoising (Kim et al., 2024; Zeng et al., 2024; Ryou et al., 2024), deblurring (Chen et al., 2024b; Wu et al., 2024a; Mao et al., 2024), deraining(Chen et al., 2024a; Pan et al., 2023; Chen et al., 2023; Ba et al., 2022; Peng et al., 2020, 2021; He et al., 2024b), dehazing(Zhang et al., 2024e; Wang et al., 2024c; Zheng et al., 2023; Wu et al., 2023), desnowing(Guo et al., 2025; Chen et al., 2021; Zhang et al., 2021) and low-light enhancement(Ma et al., 2022; Cai et al., 2023; Hai et al., 2023; Wang et al., 2022, 2023b; Ma et al., 2024). These approaches, leveraging convolutional neural networks (CNN)-based (LeCun et al., 1989), transformer-based (Vaswani et al., 2017; Pan et al., 2025a, 2024), or diffusion-based (Ho et al., 2020) architectures, enable powerful feature learning for image restoration. However, most existing works focus on addressing a single type of degradation, which limits their applicability in real-world scenarios where images often suffer from diverse and unknown types of degradation.
Recent advancements in All-in-One image restoration frameworks focus on enhancing flexibility and efficiency by leveraging various multi-task learning techniques such as Mixture-of-Experts (MoE) and text-based or vision-based prompts to handle diverse restoration tasks. For example, AirNet(Li et al., 2022) employs contrastive learning to extract degradation representations for guiding restoration. PromptIR(Potlapalli et al., 2023) integrates end-to-end prompt generation and interaction to adaptively handle different degradation types, while OneRestore(Guo et al., 2024a) aligns text-based degradation categories with visual features as prompts. IDR(Zhang et al., 2023a) uses meta-learning to decompose degradations, enabling generalization across diverse scenarios. MoFME(Zhang et al., 2024b), on the other hand, employs an uncertainty-aware router to dynamically assign Degradation-specific features to expert modules, allowing it to specialize in multiple tasks. Similarly, AdaIR(Cui et al., 2025) achieves adaptive restoration by decoupling degradations in the frequency domain. Additionally, methods like DA-CLIP(Luo et al., 2023a) leverage the pre-trained vision-language model (VLM) for prompt extraction, but the significant computational overhead introduced by VLM limits their practical efficiency. However, these methods either limit the representation of degradation information to coarse-grained degradation categories or lack a clear mechanism for extracting degradation features, which often results in suboptimal performance. In contrast, we propose a novel All-in-One image restoration framework, DPMambaIR, which is capable of extracting fine-grained degradation information and incorporating a state-space modeling process, effectively addressing diverse degradation challenges in image restoration.
2.2. State Space Model
State Space Model (SSM), originally introduced in the 1960s for control systems(Kalman, 1960), has recently gained significant attention in computer vision due to its ability to model long-range dependencies with linear computational complexity. The first effort to incorporate SSM into vision tasks was initiated by Vision Mamba(Zhu et al., 2024), which proposed the Vision State Space Module (VSSM) to extend SSM to visual applications. Vision Mamba demonstrated superior performance compared to Vision Transformers (ViTs)(Dosovitskiy et al., 2020), while maintaining lower model complexity, thereby laying a solid foundation for subsequent advancements in the field of computer vision. The success of VSSM has inspired extensive research into the application of SSM across various vision tasks, such as object detection(Zhang et al., 2024a; Yang et al., 2025; Liu et al., 2024; Li et al., 2024c; Zhang et al., 2023b, c), semantic segmentation(Xiao et al., 2024; Huang et al., 2023; Zhang et al., 2024c), and image classification(Zhang et al., 2024d; Li et al., 2024b; Yao et al., 2024; jc5sheng2024dualmamba). Among these, MambaIR(Guo et al., 2024c) is the first approach to apply SSM to image restoration. By leveraging SSM, MambaIR achieves enhanced computational efficiency and improved global feature modeling capabilities, resulting in significant improvements in perceptual quality for restoration tasks. Furthermore, other SSM-based approaches have also emerged for various image restoration tasks(Hu and Yin, 2024; Li et al., 2024d; Guo et al., 2024b; He et al., 2024c; Sun et al., 2024b; Li et al., 2024a; Wu et al., 2024c; Xia et al., 2024; Peng et al., 2025a; Di et al., 2024; Peng et al., 2025b; Pan et al., 2025b; Qi et al., 2025; Sun et al., 2024a; Wu et al., 2024b; He et al., 2024a; Du et al., 2024; He et al., 2024d; Jiang et al., 2024; Wang et al., 2024a; Li et al., 2023; Conde et al., 2024; Ren et al., 2024; Yan et al., 2023), showcasing the strong potential and promising opportunities of SSM in this domain. However, while SSM excels at modeling long-range interactions, they inherently lack a degradation-aware mechanism, which is crucial for tasks that involve handling diverse degradations. This limitation makes it difficult to balance differences between multiple tasks during modeling. This limitation hinders the performance of the model in All-in-One image restoration tasks.
3. Method
3.1. Problem Definition
All-in-One Image Restoration aims to address a wide range of image degradation types within a unified framework. Specifically, given a degraded image $\boldsymbol{I}_{D}\in\mathbb{R}^{3\times H\times W}$, where $D$ denotes the unknown degradation process, our goal is to restore the clean image $\boldsymbol{\hat{O}}\in\mathbb{R}^{3\times H\times W}$ using a unified model $\mathcal{F}$. This can be formulated as:
(1) $\boldsymbol{\hat{O}}=\mathcal{F}(\boldsymbol{I}_{D}).$
where $f$ is designed to generalize across various degradation types, such as noise, blur, compression artifacts, and weather effects, without explicit knowledge of $D$. The degradation process $D$ can be mathematically represented as:
(2) $\boldsymbol{I}_{D}=D(\boldsymbol{I}_{C})+\boldsymbol{\eta}.$
Here, $\boldsymbol{I}_{C}\in\mathbb{R}^{3\times H\times W}$ denotes the original clean image, $D$ represents the degradation operator (e.g., blur kernel, down-sampling, or quality compression), and $\boldsymbol{\eta}$ signifies additive noise. The main challenge arises from the fact that $D$ is usually unknown and varies across tasks, making it difficult to design a universal model capable of effectively handling all possible degradations.
3.2. Traditional State Space Model
In Image Restoration, Convolutional Neural Networks (CNN)(LeCun et al., 2015) and Transformer(Vaswani et al., 2017) have been widely adopted due to their strong feature extraction and representation learning capabilities. CNN excels at capturing local patterns, making them suitable for tasks requiring high spatial resolution and fine details. However, their limited receptive field often hinders their ability to model long-range dependencies, which are critical for handling global degradations. On the other hand, Transformer leverages self-attention mechanisms to capture global context effectively. Despite their success, Transformer suffers from quadratic complexity with respect to the input size, leading to high computational costs and memory consumption, especially for high-resolution images. To address these limitations, researchers have turned to Mamba, which leverages State Space Model (SSM) to achieve linear complexity and long-range dependency modeling, making it well-suited for Image Restoration tasks.
Let’s briefly review the traditional State Space Model (SSM), which maps sequence input $x(t)$ to output $y(t)$ through an implicit latent state $h(t)\in\mathbb{R}^{N}$ and can be represented as a linear ordinary differential equation:
(3) $\dot{h}(t)$ $=Ah(t)+Bx(t),$
$y(t)$ $=Ch(t)+Dx(t).$
where $N$ is the state size, $A\in\mathbb{R}^{N\times N}$, $B\in\mathbb{R}^{N\times 1}$, $C\in\mathbb{R}^{1\times N}$, and $D\in\mathbb{R}$. Discretized using a zero-order hold as:
(4) $\overline{A}$ $=\exp(\Delta A),$
$\overline{B}$ $=(\Delta A)^{-1}(\exp(\Delta A)-I)\Delta B.$
After the discretization, the discretized version of Eq. (3) with step size $\Delta$ can be rewritten as:
(5) $h_{i}$ $=\overline{A}h_{i-1}+\overline{B}x_{i},$
$y_{i}$ $=Ch_{i}+Dx_{i}.$
The advent of State Space Model has garnered significant attention for its ability to perform global modeling with linear complexity in image restoration. Existing Mamba-based methods(Zhu et al., 2024; Guo et al., 2024c, b) typically adopt specific scanning strategies to unfold 2D images into 1D sequences, enabling sequential causal modeling using Mamba. However, in All-in-One image restoration tasks, traditional SSM still faces two key limitations: (a) Limited Global Context Utilization: During the sequential causal modeling process, the $i$-th pixel can only access the previous $i-1$ pixels in the sequence, preventing effective utilization of global contextual information. To address this limitation, existing methods often employ parallel multi-directional scanning to expand the receptive field. However, this inevitably increases computational complexity. Furthermore, experiments in existing work(Guo et al., 2024c) reveal that parallel multi-directional scanning contributes minimally to performance improvements in low-level vision tasks. (b) Lack of Effective Degradation-Aware Mechanisms: Traditional SSM lack mechanisms to effectively perceive degradation information, making it difficult to balance differences between multiple tasks during modeling. This limitation hinders the performance of the model in All-in-One image restoration tasks.
3.3. DPMambaIR
We propose DPMambaIR, a degradation-aware framework for All-in-One image restoration. As illustrated in Figure 2, we first establish a Mamba-based image restoration baseline using an asymmetric encoder-decoder design within a U-shaped architecture. The restoration process begins with a $3\times 3$ convolutional layer, which extracts shallow features $\boldsymbol{F}\in\mathbb{R}^{H\times W\times C}$ from the degraded image. These features then pass through three encoding stages, each followed by downsampling. During each encoding stage, the resolution of the feature map is halved while the number of channels is doubled. Subsequently, the features pass through three decoding stages, where each stage involves upsampling and concatenating the output of the decoder with the corresponding output of the encoder to form the input for the next decoding stage. Finally, after a refinement stage, the clean image is restored through a shortcut connection that combines the input degraded image with the output of the decoder.
To address the challenges posed by unknown degradation types and levels in All-in-One image restoration tasks, we propose a Degradation Extractor pre-trained using degraded image reconstruction. This model compresses degradation type and severity information into a compact embedding, enabling the model to transition from blind to non-blind image restoration through finer-grained degradation representation. To further optimize the utilization of the extracted degradation embedding, we revisit the selective state-space modeling process. Specifically, we propose a Degradation-Aware Prompt State Space Model, which guides the image restoration process by compressing historical information and supplementing global degradation information.
3.4. Pretrained Degradation Extractor
Figure 3. (a) Degradation Modeling Process of Real-World Images, (b) The Architecture of Degradation Extractor.
Figure 4. The Architecture of (a) Degradation-aware Prompt Selective Scan (DPSS) and (b) High-frequency Enhancement Block (HEB), with Three Inputs $F_{1}$, $F_{2}$, and Degradation Embedding $D$.
To better extract degradation information from images, we revisit the degradation modeling process of real-world images, which spans from image capture and imaging to transmission, as illustrated in Figure  3 (a). This process typically involves four common types of degradations: (1) poor illumination conditions, (2) relative motion between the target objects and the camera, (3) artifacts introduced by transmission media in the optical path, and (4) compression during storage and transportation. We extend these four types of degradations to a more general framework and propose a unified degradation imaging model. The specific formulation is defined as follows:
(6) $I(x)=Q\big{(}(\alpha(x)J(x)+\beta(x)\gamma(x))\otimes K\big{)}.$
where $J(x)$ represents the clean image, $\alpha(x)$ models the luminance adjustment, such as the effects of insufficient lighting in low-light or adverse weather conditions (e.g., rain or fog), and $\gamma(x)$ accounts for artifacts caused by occlusions from aerosols, raindrops, or snow particles in the transmission medium. $\beta(x)$ modulates $\gamma(x)$ based on varying spatial properties, $K$ represents the blur introduced by relative motion between the target and the camera, and $Q$ denotes the compression applied during storage and transmission.
Based on the aforementioned degradation modeling process, we design a universal degradation extractor, as illustrated in Figure 3 (b). Specifically, we introduce central difference convolution(CDC)(Wang et al., 2023a; Peng et al., 2024a, b) to extract edge and texture degradation cues related to illumination and blur. In parallel, we employ three groups of CDC with different scales to capture artifacts of varying sizes, such as visual artifacts and snowflakes. Additionally, we integrate a traditional convolution branch to learn brightness-related representations.
Furthermore, to train the degradation extractor with fine-grained degradation encoding capability, we propose a pretraining method for the degradation extractor based on degraded image reconstruction. This method leverages a hybrid dataset consisting of paired data with diverse types of degradation. The core of this process is training the degradation extractor $\mathcal{E}$ to extract fine-grained degradation information from the degraded image $\boldsymbol{I}_{D}$ and distill it into a compact degradation embedding $D$ of size $1\times 512$, formulated as $D=\mathcal{E}(\boldsymbol{I}_{D})$. Subsequently, the reconstruction unit $\mathcal{R}$ utilizes $D$ and the clean image $\boldsymbol{O}$ to reconstruct the degraded image $\boldsymbol{I}_{D}$, formulated as $\boldsymbol{\hat{I}}_{D}=\mathcal{R}(\boldsymbol{O},D)$.
We further introduce a dual-component loss function $L$, which combines the $L_{1}$ norm and the Learned Perceptual Image Patch Similarity (LPIPS) metric. The combined loss function is expressed as $L(\boldsymbol{I}_{D},\boldsymbol{\hat{I}}_{D})=L_{1}+L_{LPIPS}$, striking a careful balance between reconstruction accuracy and perceptual quality.
3.5. Degradation-aware Prompt State Space Model
To address these limitation mentioned in 3.2, we propose a novel Degradation-aware Prompt State Space Model to fully unlock the potential of SSM in All-in-One image restoration. Specifically, we propose a novel Degradation-aware Prompt Selective Scan, which incorporates degradation information into the State Space Model. This is achieved by leveraging a Degradation Embedding extracted from a pre-trained Degradation Extractor to provide supplementary degradation-related contextual information that the traditional SSM cannot capture. In addition, we modulate the parameter matrices in the state-space equations to establish a degradation-aware mechanism. This mechanism enables SSM to adapt to the imbalances caused by the inherent differences among sub-tasks in All-in-One image restoration. To further mitigate the issue of high-frequency information loss, which arises from balancing the differences among sub-tasks during training, we design a simple yet effective High-frequency Enhancement Block (HEB). This component ensures the preservation of high-frequency details and significantly enhances the overall performance of the model.
Degradation-aware Prompt Selective Scan. We revisit the traditional SSM, whose parameter settings are shown in Equation (3). Matrices $A$, $B$, $C$, and $D$ control the SSM output. The state transition matrix $A$ compresses historical information, while the input matrix $B$ maps input signals to the state space, determining their influence on historical states. The output matrix $C$ maps historical states to the observable space, reflecting their impact on the output. Matrix $D$ provides a direct pathway from input to output, similar to a residual connection. In traditional SSM, these matrices and $\Delta$ are derived from:
(7) $A=init(),D=init(),$
$B=Linear(x),C=Linear(x),\Delta=Linear(x).$
where $init()$ is an initialization method that does not depend on the input $x$. This way may lead to two key issues: (a) Traditional State Space Model struggles to effectively leverage global contextual information. (b) The lack of a proper degradation-aware mechanism makes it challenging to balance the differences among tasks in multi-task learning.
To address these issues, we first utilize a pre-trained degradation extractor $\mathcal{E}$ to extract the degradation embedding $D$ from the input image $I_{D}$. Subsequently, $D$ is used to modulate $\Delta$, the input matrix $B$, and the output matrix $C$ of the SSM. This introduces a degradation-aware mechanism in three aspects: historical information compression, input information processing, and global context awareness, as illustrated in Figure 4 (a). The specific method is formulated as follows:
(8) $D$ $=\mathcal{E}(I_{D}),$
$\alpha_{\Delta}$ $=\mathcal{M}_{\Delta}(D),\,\alpha_{B}=\mathcal{M}_{B}(D),\,\alpha_{C}=\mathcal{M}_{C}(D),$
$\Delta_{dp}$ $=\alpha_{\Delta}\cdot\Delta,\,B_{dp}=\alpha_{B}\cdot B,\,C_{dp}=\alpha_{C}\cdot C.$
Here, $\mathcal{M}_{\Delta}$, $\mathcal{M}_{B}$, and $\mathcal{M}_{C}$ are combinations of linear layers used to map the degradation embedding $\mathbf{e}$ into feature spaces, forming modulation vectors $\alpha_{\Delta}$, $\alpha_{B}$, and $\alpha_{C}$ for $\Delta$, $B$, and $C$, respectively. $\Delta_{dp}$, $B_{dp}$, and $C_{dp}$ are the newly modulated SSM parameters. Finally, the state-space function formulation of the Degradation-aware State Space Model can be expressed as follows:
(9) $\overline{A}_{dp}$ $=\exp(\Delta_{dp}A),$
$\overline{B}_{dp}$ $=(\Delta_{dp}A)^{-1}(\exp(A)-I)\cdot\Delta_{dp}B_{dp},$
$h_{i}$ $=\overline{A}_{dp_{i}}h_{i-1}+\overline{B}_{dp_{i}}x_{i},$
$y_{i}$ $=C_{dp}h_{i}+Dx_{i}.$
Through this approach, the transformation matrix $A$ and input matrix $B$ gain degradation-aware capabilities, integrating degradation cues of various types and levels during historical information compression and input signal processing, making the model more task-friendly. The output matrix $C$ controls how historical information affects the output. In traditional SSM, the $i$-th pixel can only observe the $i-1$ preceding pixels, leading to a limited global receptive field. While existing methods address this with parallel multi-directional scanning, this increases computational costs and offers limited performance gains in low-level vision tasks (Guo et al., 2024c). Instead, we retain single-directional scanning and enhance global degradation cues via $C$, achieving efficient global awareness with lower computational overhead.
High-frequency Enhancement Block.
Figure 5. We compare frequency-domain differences between clear and restored images under three configurations. The All-in-One setting shows a notable high-frequency gap compared to the Degradation-specific setting, which is significantly reduced by adding the High-frequency Enhancement Block (HEB), as shown by the blue line.
In All-in-One image restoration tasks, models tend to prioritize low-frequency restoration over high-frequency details, such as edges and textures, leading to their loss, as shown in Figure 5. To address this, we propose an efficient high-frequency enhancement block with minimal computational overhead. For input features $F$, we first extract channel-wise low-frequency features using global average pooling, then subtract them from $F$ to obtain high-frequency features, as shown in Figure 4(b). Finally, the high-frequency features are added back to enhance $F$, formulated as:
(10) ${F}_{enhanced}={F}+\alpha\cdot\big{(}{F}-\mathcal{G}({F})\big{)}.$
Here, $\alpha$ represents a learnable parameter that controls the weight of high-frequency information during the enhancement process. $\mathcal{G}$ represents the global average pooling operation. ${F}_{enhanced}$ denotes the features after high-frequency enhancement.
3.6. Loss Function
We employ the commonly used L1 loss $\mathcal{L}_{1}$ and L2 loss $\mathcal{L}_{2}$ in low-level vision tasks. Additionally, following previous works(Zamir et al., 2022; Cui et al., 2023), we incorporate a frequency-domain loss $\mathcal{L}_{fft}$ for training. The total loss is defined as:
(11) $\mathcal{L}_{total}=\lambda_{1}\cdot\mathcal{L}_{1}(O,\hat{O})+\lambda_{2}\cdot\mathcal{L}_{2}(O,\hat{O})+\lambda_{3}\cdot\mathcal{L}_{fft}(O,\hat{O}).$
where $O$ and $\hat{O}$ denote the model output and the ground truth, respectively. The parameters $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are balancing factors, which we set to 1, 0.5 and 0.001, respectively.
4. Experiment And Analysis
We evaluate our method on both All-in-one image restoration tasks and Degradation-specific image restoration tasks. In the All-in-One Image Restoration setting, we train a single model on a dataset containing seven types of degradation. For Degradation-specific image restoration, we train separately on datasets corresponding to each type of degradation.
4.1. Experimental Settings
Training Details. Building on previous works(Zamir et al., 2022; Potlapalli et al., 2023; Guo et al., 2024c), we develop a Mamba-based baseline for image restoration. Our method is trained using the AdamW(Kingma and Ba, 2014) optimizer with parameters $\beta_{1}=0.9$, $\beta_{2}=0.999$, and a weight decay of $1e^{-4}$ for 300,000 iterations. The initial learning rate of $3e^{-4}$ is gradually reduced to $1e^{-6}$ through cosine annealing. Following(Zamir et al., 2022), we employ a progressive training strategy, cropping image patches of size $192\times 192$. Horizontal and vertical flips are applied for data augmentation. Our experiments are conducted on a single NVIDIA RTX A100 GPU and implemented using the PyTorch platform.
Datasets. For All-in-One image restoration, we collect a large-scale mixed dataset that includes seven classic and popular degradation types: fog, rain, snow, low light, blur, JPEG compression artifacts, and noise. Additionally, we evaluate performance on Degradation-specific tasks, using the LOL dataset(Wei et al., 2018) for low-light image enhancement, the GoPro dataset(Nah et al., 2017a) for image deblurring, the Rain100H dataset(Yang et al., 2017) for image deraining, and the RESIDE6K dataset(Qin et al., 2020) for image dehazing.
Table 1. Quantitative Comparison on Seven Datasets for All-in-One Image Restoration Task with State-of-the-Art Methods. The Best and Second-Best Results Are in Bold and Underline.
Method Lowlight Snow Rain Fog Jpeg Noise Blur Average PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM MPRNet 19.71 0.659 30.94 0.938 26.27 0.912 25.63 0.947 30.51 0.938 28.03 0.897 26.07 0.869 26.74 0.880 MIRNet 19.81 0.668 31.04 0.941 25.77 0.907 22.26 0.882 30.54 0.939 28.14 0.900 26.32 0.873 26.27 0.873 NAFNet 19.49 0.650 30.33 0.935 26.04 0.906 25.63 0.946 30.20 0.936 27.80 0.895 26.13 0.870 26.52 0.877 Restormer 19.99 0.670 31.67 0.944 26.91 0.919 26.18 0.953 30.56 0.939 28.17 0.890 26.62 0.880 27.16 0.886 PromptIR 19.90 0.669 31.52 0.943 26.80 0.918 26.13 0.945 30.57 0.939 28.18 0.901 26.36 0.874 27.07 0.884 IDR 19.68 0.663 31.16 0.941 26.39 0.914 26.11 0.948 30.51 0.939 28.10 0.899 26.29 0.873 26.89 0.883 OneRestore 19.77 0.657 29.95 0.931 25.37 0.900 27.74 0.962 30.10 0.934 27.65 0.890 25.90 0.866 26.64 0.877 AdaIR 19.99 0.672 31.94 0.923 27.01 0.922 26.61 0.952 30.53 0.939 28.19 0.900 26.61 0.879 27.26 0.887 DPMambaIR 20.04 0.680 32.36 0.949 27.07 0.926 28.13 0.963 30.57 0.939 28.25 0.902 27.43 0.894 27.69 0.893
Table 2. Quantitative Comparison on Deraining, Low-light Image Enhancement(LLIE), Deblurring and Dehazing for Degradation-Specific Image Restoration Task with State-of-the-Art Methods. The Best and Second-Best Results Are in Bold and Underline.
Deraining PSNR SSIM JORDER 26.25 0.835 PReNet 29.46 0.899 MPRNet 30.41 0.891 MAXIM 30.81 0.903 Restormer 31.46 0.904 DPMambaIR 32.30 0.967
LLIE PSNR SSIM EnlightenGAN 17.61 0.653 MIRNet 24.14 0.830 URetinex-Net 19.84 0.824 MAXIM 23.43 0.863 IAT 23.38 0.809 DPMambaIR 24.20 0.852
Deblurring PSNR SSIM DeepDeBlur 29.08 0.913 DeBlurGAN 28.70 0.858 DeBlurGANV2 29.55 0.934 MT-RNN 31.15 0.945 IR-SDE 30.70 0.901 DPMambaIR 31.42 0.948
Dehazing PSNR SSIM GCANet 26.59 0.935 GridDehazeNet 25.86 0.944 DeHazeFormer 30.29 0.964 MAXIM 29.12 0.932 DA-CLIP 30.16 0.936 DPMambaIR 30.89 0.977
Figure 6. Qualitative comparison on Deblurring, Deraining and Dehazing for All-in-One Image Restoration Task.
Evaluation Metrics. Following previous works(Zamir et al., 2022; Guo et al., 2024c), we use peak signal-to-noise ratio (PSNR) and structural similarity index measure (SSIM) to quantitatively compare different methods on the RGB channels.
Comparisons with State-of-the-art Methods. For All-in-One image restoration, we compare our method against eight approaches: MIRNet (Zamir et al., 2020), NAFNet (Chen et al., 2022), MPRNet (Zamir et al., 2021), Restormer (Zamir et al., 2022), PromptIR (Potlapalli et al., 2023), IDR (Zhang et al., 2023a), OneRestore (Guo et al., 2024a), and AdaIR (Cui et al., 2025). Among these, MIRNet, NAFNet, MPRNet, and Restormer are general-purpose image restoration models, while the latter four methods are specifically designed for All-in-One image restoration. Specifically, PromptIR employs visual prompts to guide restoration, OneRestore leverages a pre-trained degradation category encoder with cross-attention mechanisms, and AdaIR utilizes frequency domain information for adaptive restoration. For Degradation-specific image restoration tasks, we compare our method with different state-of-the-art approaches tailored to each domain. The details are as follows: (1) For deraining, we compare against JORDER (Yang et al., 2019), PReNet (Ren et al., 2019), MPRNet (Zamir et al., 2021), MAXIM (Tu et al., 2022), and Restormer (Zamir et al., 2022). (2) For low-light image enhancement, we compare against EnlightenGAN (Jiang et al., 2021), MIRNet (Zamir et al., 2020), Uretinex-Net (Wu et al., 2022), MAXIM (Tu et al., 2022), and IAT (Cui et al., 2022). (3) For deblurring, we compare against DeepDeblur (Nah et al., 2017b), DeblurGAN (Kupyn et al., 2017), DeblurGANv2 (Kupyn et al., 2019), MT-RNN (Park et al., 2020), and IR-SDE (Luo et al., 2023b). (4) For dehazing, we compare against GCANet (Chen et al., 2018), GridDehazeNet (Liu et al., 2019), DeHazeFormer (Song et al., 2023), MAXIM (Tu et al., 2022), and DA-CLIP (Luo et al., 2023a).
4.2. Quantitative Results
All-in-one Image Restoration. We compare our method against five state-of-the-art image restoration approaches, including both CNN-based and transformer-based methods. As shown in Table 1, for the All-in-One image restoration task, our method achieves the best PSNR and SSIM on the aforementioned mixed datasets compared to existing methods. Specifically, our method achieves a PSNR of 27.69dB and an SSIM of 0.893, outperforming AdaIR by 0.43dB and 0.011, respectively. We achieve the best performance across all seven sub-tasks. Experimental results indicate that multi-task learning often encounters challenges in balancing performance across different tasks. For instance, AdaIR demonstrates strong performance in deraining but performs poorly in dehazing. On the other hand, OneRestore excels in dehazing but delivers suboptimal results in tasks such as deraining and deblurring. In contrast, our method effectively tackles the balance issue in multi-task learning, achieving consistently optimal performance across all sub-tasks. These findings highlight the effectiveness and superiority of our proposed method.
Degradation-Specific Image Restoration. We evaluate our method on four Degradation-specific tasks, consistently achieving superior performance across multiple datasets compared to state-of-the-art CNN-based, transformer-based, and diffusion-based approaches (Table 2). On Rain100H (Yang et al., 2017), our method achieves a PSNR of 32.30dB and SSIM of 0.967, surpassing Restormer by 0.84dB and 0.063, respectively. On the LOL dataset (Wei et al., 2018), it achieves 24.20dB PSNR and 0.852 SSIM, outperforming MIRNet by 0.06dB and 0.022. For deblurring on GoPro (Nah et al., 2017a), it attains 31.42dB PSNR and 0.948 SSIM, exceeding MT-RNN by 0.27dB and 0.003. On RESIDE6K (Qin et al., 2020), it achieves 30.89dB PSNR and 0.973 SSIM, outperforming DehazeFormer by 0.60dB and 0.009. These results highlight the versatility and effectiveness of our method in addressing diverse degradation restoration tasks.
4.3. Qualitative Results
Figure 6 presents the visual results of deblurring, deraining, and dehazing under the All-in-One Image Restoration task setting. Our method demonstrates superior visual performance compared to other approaches. Specifically, in deblurring, our method restores sharper edge details; in deraining, it more effectively removes rain streaks while recovering background information; and in dehazing, it focuses on regions often overlooked by other methods, resulting in more comprehensive dehazing performance. Additional visual comparisons under both the All-in-One and Degradation-specific settings are provided in the Appendix.
4.4. Ablation Study
Table 3 presents the results of our ablation study on DPMambaIR for the All-in-One image restoration task, illustrating the impact of removing the degradation prompt, different modulation modes, and the High-frequency Enhancement Block (HEB) from the model. For the baseline model without the Degradation-Aware Prompt Selective Scan and HEB, the PSNR and SSIM are only 26.96 dB and 0.884, respectively. Introducing degradation-awareness via prompts applied to Matrices $A$ and $B$ improves PSNR and SSIM by 0.64 dB and 0.008, respectively, by enabling degradation-aware mechanisms in historical information compression and input feature processing. Adding Prompt $C$, which supplements global degradation information, further increases PSNR by 0.07 dB. The inclusion of HEB significantly boosts performance, preserving high-frequency details. The complete model achieves 27.69 dB PSNR and 0.893 SSIM, demonstrating the importance of all components for optimal performance. Table 4 presents the results for degradation extraction and utilization methods. For degradation extraction, we compare three approaches: explicit category labels, a pre-trained degradation extractor from OneRestore (Guo et al., 2024a), and our fine-grained degradation representation via image reconstruction. The OneRestore extractor, designed for classification tasks, performs worse than explicit category labels. In contrast, our method achieves a 0.11 dB PSNR improvement over non-blind restoration with category labels by leveraging fine-grained degradation representations. For degradation utilization, we compare concatenation and attention, with our method outperforming both, demonstrating its effectiveness and superiority in extracting and utilizing degradation information.
For the ablation study on global information supplementation, we compare our method with Mamba’s dual-sequence scanning scheme. As shown in Table 5, our approach not only achieves superior performance compared to the dual-sequence scanning method but also demonstrates a significant reduction in both the Parameters and FLOPs.
Table 3. Ablation Studies on our proposed core Modules.
[…]
Table 4. Ablation Study on Different Degradation Extraction and Utilization Methods.
Comparison Projects Method PSNR SSIM Degradation Extraction Classes 27.58 0.892 OneRestore* 27.56 0.891 Degradation Insertion Concat 26.96 0.882 Attention 27.39 0.882 Ours 27.69 0.893
Table 5. Ablation Study on Different Methods for Global Information Supplementation.
Method PSNR SSIM Params(M) GFLOPs Bi-direction Scan 27.66 0.893 31.43 244.3 Prompt Matrix C 27.69 0.893 29.35 153.3
5. Conclusion
In this paper, we proposed DPMambaIR, a novel All-in-One framework for image restoration that balances high restoration quality and computational efficiency across diverse degradation types. DPMambaIR introduces a Degradation-Aware Prompt State Space Model (DP-SSM), which dynamically integrates fine-grained degradation features into the state-space modeling process. This design enables adaptive handling of complex degradation scenarios while maintaining global degradation awareness. Furthermore, the High-Frequency Enhancement Block (HEB) reduces high-frequency detail loss caused by task competition, ensuring superior edge and texture restoration. This study underscores the importance of fine-grained degradation modeling and dynamic feature modulation in advancing All-in-One image restoration frameworks. Extensive experiments on a mixed dataset containing seven degradation types show that DPMambaIR achieves the best performance, establishing it as a promising approach for unified image restoration.