### v2

# DPMambaIR: All-in-One Image Restoration via Degradation-Aware Prompt State Space Model

Abstract
All-in-One image restoration aims to address multiple image degradation problems using a single model, offering a more practical and versatile solution compared to designing dedicated models for each degradation type. Existing approaches typically rely on Degradation-specific models or coarse-grained degradation prompts to guide image restoration. However, they lack fine-grained modeling of degradation information and face limitations in balancing multi-task conflicts. To overcome these limitations, we propose DPMambaIR, a novel All-in-One image restoration framework that introduces a fine-grained degradation extractor and a Degradation-Aware Prompt State Space Model (DP-SSM). The DP-SSM leverages the fine-grained degradation features captured by the extractor as dynamic prompts, which are then incorporated into the state space modeling process. This enhances the model’s adaptability to diverse degradation types, while a complementary High-Frequency Enhancement Block (HEB) recovers local high-frequency details. Extensive experiments on a mixed dataset containing seven degradation types show that DPMambaIR achieves the best performance, with 27.69dB and 0.893 in PSNR and SSIM, respectively. These results highlight the potential and superiority of DPMambaIR as a unified solution for All-in-One image restoration.
I Introduction
Image restoration is a fundamental task in computer vision, aiming to recover high-quality images from degraded inputs. Recent advances in deep learning have significantly improved performance in Degradation-specific restoration tasks, such as image denoising [1, 2], deblurring [3], deraining[4], dehazing[5, 6] and low-light enhancement[7]. However, traditional methods typically rely on Degradation-specific models explicitly tailored to individual degradation types. These approaches often require prior knowledge of the degradation type, which limits their practicality in real-world scenarios, such as autonomous driving and nighttime surveillance, where diverse and unknown degradations frequently coexist.
Figure 1: Comparison with SOTA All-in-One Image Restoration Methods on mixed Dataset. Our DPMambaIR achieves consistently better performance.
To address this limitation, recent studies explore All-in-One image restoration frameworks that unify multiple restoration tasks into a single model. Existing methods can be broadly categorized into Mixture-of-Experts (MoE)-based and Prompt-based approaches. MoE-based methods, such as MoFME [8], dynamically route tasks to specialized experts using an uncertainty-aware router, enabling effective learning of diverse expert functions. However, these methods suffer from high computational costs, increased inference complexity as the number of experts grows, and lack explicit degradation modeling, limiting adaptability to unseen or complex degradations. Prompt-based methods, like OneRestore [9], guide restoration using degradation prompts, integrating scene descriptors via cross-attention mechanisms for flexible and adaptive strategies. While effective, these methods rely on coarse prompts that fail to capture nuanced degradation characteristics and treat prompts as static inputs, limiting their ability to adapt to varying degradation intensities or specific requirements, leading to suboptimal results.
While these methods have demonstrated success in learning multiple tasks for All-in-One image restoration, two key challenges remain unresolved. First, existing models lack a degradation extractor capable of capturing fine-grained degradation features and often rely on coarse degradation labels to guide the restoration process. For example, images with varying degrees of haze are typically assigned a single haze label, failing to capture the nuanced variations in degradation levels. This simplification leads to suboptimal restoration performance, as the model cannot adapt to the diverse characteristics of different degradations. Second, current methods struggle to balance the competing requirements of varying restoration tasks. For instance, training low-light enhancement, which focuses on restoring low-frequency contrast information, alongside deblurring, which prioritizes the recovery of fine local details. This inherent task competition makes it challenging to optimize a single network for all restoration tasks simultaneously and remains a fundamental challenge for All-in-One image restoration frameworks.
To address these challenges, we propose DPMambaIR, a novel All-in-One image restoration framework built upon a regression-based degradation extractor and a Degradation-Aware Prompt State Space Model (DP-SSM). Unlike prior approaches that use static or class-based prompts, our method employs a degradation extractor to generate prompt vectors that encode both semantic categories and degradation distribution characteristics. These prompts are then used to dynamically modulate the state transition, input, and output matrices of a state space model, enabling fine-grained degradation modeling and adaptive coordination across frequency bands. This mechanism allows our model to better manage task competition and enhance restoration quality across diverse degradations. Moreover, while State Space Model such as Mamba exhibit strong global modeling capabilities, they are inherently biased toward low-frequency restoration [10]. To better recover local high-frequency details, we introduce a lightweight High-Frequency Enhancement Block (HEB) as a supplementary component, which delivers performance gains while incurring only negligible computational overhead.
Extensive experiments on a mixed dataset containing seven degradation types demonstrate the superiority of DPMambaIR. As shown in Fig. 1, DPMambaIR achieves state-of-the-art performance in terms of PSNR and SSIM, outperforming existing All-in-One methods such as AdaIR and OneRestore. Moreover, DPMambaIR achieves competitive results across individual tasks, including deraining, low-light enhancement, deblurring, and dehazing, showcasing its potential as an effective and robust All-in-One image restoration solution.
The contributions of this paper are summarized as follows:
  • $\bullet$
    We propose a novel degradation extractor capable of capturing fine-grained degradation features from complex degraded images. These features provide critical priors to improve the performance of All-in-One image restoration.
  • $\bullet$
    We design a Degradation-aware Prompt State Space Model (DP-SSM), which enhances the long-range modeling capability of Mamba by integrating degradation priors at multiple structural levels.
  • $\bullet$
    Extensive experiments on a mixed dataset containing seven types of degradation demonstrate that DPMambaIR achieves state-of-the-art performance on PSNR and SSIM metrics, outperforming existing methods. The results validate the effectiveness and superiority of the proposed approach.
II Related Work
II-A All-in-One Image Restoration
Image restoration is a fundamental task in computer vision, aiming to recover high-quality images from degraded inputs. Traditional methods, such as Dark Channel Prior [11] and Color Line Prior [12], rely on handcrafted priors to model degradation but struggle with complex scenarios due to predefined assumptions. The rise of deep learning [13] has significantly advanced tasks like such as image super-resolution[14, 15, 16], image denoising [1, 2], deblurring [3], deraining[4], dehazing[5], desnowing[17] and low-light enhancement[7]. These approaches, leveraging convolutional neural networks (CNN)-based [18], transformer-based [19], or diffusion-based [20] architectures, enable powerful feature learning for image restoration. However, most existing works focus on addressing a single type of degradation, which limits their applicability in real-world scenarios where images often suffer from diverse and unknown types of degradation.
Recent advancements in All-in-One image restoration frameworks focus on enhancing flexibility and efficiency by leveraging various multi-task learning techniques such as Mixture-of-Experts (MoE) and text-based or vision-based prompts to handle diverse restoration tasks. For example, AirNet[21] employs contrastive learning to extract degradation representations for guiding restoration. PromptIR[22] integrates end-to-end prompt generation and interaction to adaptively handle different degradation types, while OneRestore[9] aligns text-based degradation categories with visual features as prompts. IDR[23] uses meta-learning to decompose degradations, enabling generalization across diverse scenarios. MoFME[8], on the other hand, employs an uncertainty-aware router to dynamically assign Degradation-specific features to expert modules, allowing it to specialize in multiple tasks. Similarly, AdaIR[24] achieves adaptive restoration by decoupling degradations in the frequency domain. Additionally, methods like DA-CLIP[25] leverage the pre-trained vision-language model (VLM) for prompt extraction, but the significant computational overhead introduced by VLM limits their practical efficiency.
Despite their effectiveness, most prompt-based methods rely on coarse-grained class labels or static degradation tokens, which limits their ability to model continuous degradation distributions or adapt to image-specific degradation characteristics. In contrast, our method extracts fine-grained, regression-level degradation embeddings that encode both semantic and statistical degradation features. These embeddings are used to dynamically modulate the state-space parameters of the restoration backbone, enabling degradation-aware processing throughout the network. This facilitates better coordination across restoration tasks and reduces competition between frequency components.
Figure 2: The overall framework of DPMambaIR consists of a shallow feature extractor and multiple DPSS blocks, with an upsampler implemented by pixel shuffle and a downsampler by pixel unshuffle. Additionally, a degradation extractor is included to extract degradation embeddings, which are then inserted into the DPSS blocks.
II-B State Space Model based Image Restoration
In Image Restoration, Convolutional Neural Networks (CNN) [13] and Transformer [19] have been widely adopted due to their strong feature extraction and representation learning capabilities. CNN excels at capturing local patterns, making them suitable for tasks requiring high spatial resolution and fine details. However, their limited receptive field often hinders their ability to model long-range dependencies, which are critical for handling global degradations. On the other hand, Transformer leverages self-attention mechanisms to capture global context effectively. Despite their success, Transformers suffer from quadratic complexity relative to the input size, making them computationally and memory-intensive.
To address these limitations, researchers have turned to the State Space Model (SSM)[26], which provides an efficient framework for modeling long-range dependencies with linear computational complexity. Formally, SSM maps an input sequence $x(t)$ to an output sequence $y(t)$ through an implicit latent state $h(t)\in\mathbb{R}^{N}$, represented by a linear ordinary differential equation:
$h^{\prime}(t)$ $=Ah(t)+Bx(t),$ (1)
$y(t)$ $=Ch(t)+Dx(t).$
where $A\in\mathbb{R}^{N\times N}$, $B\in\mathbb{R}^{N\times 1}$, $C\in\mathbb{R}^{1\times N}$, and $D\in\mathbb{R}$. After discretization with step size $\Delta$, the dynamic equation can be expressed as:
$h_{i}$ $=\overline{A}h_{i-1}+\overline{B}x_{i},$ (2)
$y_{i}$ $=Ch_{i}+Dx_{i}.$
where $\overline{A}=\exp(\Delta A)$ and $\overline{B}=(\Delta A)^{-1}(\exp(\Delta A)-I)\Delta B$.
The introduction of Vision Mamba[27] marked the first effort to incorporate SSM into vision tasks. By proposing the Vision State Space Module (VSSM), it extended SSM to visual applications and demonstrated superior performance over Vision Transformer (ViT)[28], while maintaining lower complexity. This success inspired extensive research into SSM across diverse tasks such as object detection[29], semantic segmentation[30], and image classification[31].
In the context of Image Restoration, MambaIR[32] pioneered the integration of SSM, leveraging its linear complexity and long-range modeling capability to achieve significant improvements in both computational efficiency and perceptual quality. This was followed by several variants targeting specific restoration tasks[33, 34, 35], collectively showcasing the strong potential of SSM in this domain. Existing approaches typically unfold 2D images into 1D sequences and apply sequential causal modeling. However, this design still faces two limitations: (a) Limited Global Context Utilization: The $i$-th pixel can only access the previous i1i-1 pixels in the sequence, restricting the effective use of global context. Parallel multi-directional scanning has been explored to enlarge the receptive field, but it increases computational cost while providing limited gains in low-level vision tasks[32]. (b) Lack of Effective Degradation-Aware Mechanisms: Current SSM-based methods primarily target single-task restoration. Their transition and output mappings are fixed or solely input-driven, without leveraging external degradation priors. As a result, they cannot dynamically adapt to diverse degradation types or levels, leading to suboptimal performance in All-in-One image restoration task.
These limitations motivate the introduction of a degradation-aware modulation mechanism into the SSM pipeline. In this work, we propose a Degradation-aware Prompt State Space Model (DP-SSM), where degradation embeddings dynamically modulate the state-space transition and output matrices. This design enables a fine-grained, task-aware mechanism that balances global modeling with degradation-specific guidance, enhancing the flexibility and effectiveness of SSM for complex restoration scenarios.
III Method
III-A Problem Definition
All-in-One Image Restoration aims to address a wide range of image degradation types within a unified framework. Specifically, given a degraded image $\boldsymbol{I}_{D}\in\mathbb{R}^{3\times H\times W}$, where $D$ denotes the degradation process, our goal is to restore the clean image $\boldsymbol{\hat{O}}\in\mathbb{R}^{3\times H\times W}$ using a unified model $\mathcal{F}$. This can be formulated as:
$\boldsymbol{\hat{O}}=\mathcal{F}(\boldsymbol{I}_{D}).$ (3)
where $\mathcal{F}$ is designed to generalize across various degradation types, such as noise, blur, compression artifacts, and weather effects, without explicit knowledge of $Deg$. The degradation process $Deg$ can be mathematically represented as:
$\boldsymbol{I}_{D}=Deg(\boldsymbol{I}_{C})+\boldsymbol{\eta}.$ (4)
Here, $\boldsymbol{I}_{C}\in\mathbb{R}^{3\times H\times W}$ denotes the original clean image, $Deg$ represents the degradation operator (e.g., blur kernel, brightness, or quality compression), and $\boldsymbol{\eta}$ signifies additive noise. The main challenge arises from the fact that $Deg$ is usually unknown and varies across tasks, making it difficult to design a universal model capable of effectively handling all possible degradations.
Unlike degradation-specific restoration, the All-in-One setting requires the model to generalize across a broad spectrum of degradation types without prior knowledge of degradation parameters. This introduces two main challenges: (a) how to represent fine-grained degradation information in a unified way; and (b) how to dynamically adapt the model behavior based on such information during inference.
Figure 3: (a) Real-World Image Degradation Modeling Process. (b) Architecture details of the Large version degradation extractor and reconstructor. (c) Architecture details of the normal version degradation extractor and reconstructor.
III-B DPMambaIR
We propose DPMambaIR, a degradation-aware framework for All-in-One image restoration. As illustrated in Fig. 2, we first establish a Mamba-based image restoration baseline using an asymmetric encoder-decoder design within a U-shaped architecture. The restoration process begins with a $3\times 3$ convolutional layer, which extracts shallow features $\boldsymbol{F}\in\mathbb{R}^{H\times W\times C}$ from the degraded image. These features then pass through three encoding stages, each followed by downsampling. During each encoding stage, the resolution of the feature map is halved while the number of channels is doubled. Subsequently, the features pass through three decoding stages, where each stage involves upsampling and concatenating the output of the decoder with the corresponding output of the encoder to form the input for the next decoding stage. Finally, after a refinement stage, the clean image is restored through a shortcut connection that combines the input degraded image with the output of the decoder.
To address the challenges posed by unknown degradation types and levels in All-in-One image restoration tasks, we propose a Degradation Extractor pre-trained using degraded image reconstruction. This model compresses degradation type and severity information into a compact embedding, enabling the model to transition from blind to non-blind image restoration through finer-grained degradation representation. To further optimize the utilization of the extracted degradation embedding, we revisit the selective state-space modeling process. Specifically, we propose a Degradation-Aware Prompt State Space Model, which guides the image restoration process by compressing historical information and supplementing global degradation information.
III-C Degradation Extractor
To better extract degradation information from images, we revisit the degradation modeling process of real-world images, which spans from image capture and imaging to transmission, as illustrated in Fig. 3 (a). This process typically involves four common types of degradations: (1) poor illumination conditions, (2) relative motion between the target objects and the camera, (3) artifacts introduced by transmission media in the optical path, and (4) compression during storage and transportation. We unify these scenarios into a generalized degradation formation model:
$I(x)=Q\big((\alpha(x)J(x)+\beta(x)\gamma(x))\otimes K\big).$ (5)
where $J(x)$ represents the clean image, $\alpha(x)$ models the luminance adjustment, such as the effects of insufficient lighting in low-light or adverse weather conditions (e.g., rain or fog), and $\gamma(x)$ accounts for artifacts caused by occlusions from aerosols, raindrops, or snow particles in the transmission medium. $\beta(x)$ modulates $\gamma(x)$ based on varying spatial properties, $K$ represents the blur introduced by relative motion between the target and the camera, and $Q$ denotes the compression applied during storage and transmission.
Figure 4: The Architecture of (a) Degradation-aware Prompt Selective Scan (DPSS) and (b) High-frequency Enhancement Block (HEB), with Three Inputs $F_{1}$, $F_{2}$, and Degradation Embedding $\boldsymbol{E}_{d}$.
Based on the aforementioned degradation modeling process, we design a universal degradation extractor, as illustrated in Fig. 3 (b). Specifically, we introduce central difference convolution (CDC)[36, 37, 38] to extract edge and texture degradation cues related to illumination and blur. In addition, we propose a multi-scale degradation perception module, which employs four parallel CDC operators of different sizes to capture degradation features at various scales. By concatenating the outputs, the module retains complete degradation representations across all frequency components. Furthermore, we employ three groups of CDC with different scales to capture artifacts of varying sizes, such as visual artifacts and snowflakes. Additionally, we integrate a traditional convolution branch to learn brightness-related representations.
To achieve a balance between extraction performance and computational cost, we introduce a re-parameterization technique for efficient integration of degradation representations, as illustrated in Fig. 3 (c). These two implementations correspond to different variants of our method, which are presented in Table I as DPMambaIR-L and DPMambaIR, respectively.
Furthermore, to endow the degradation extractor with fine-grained degradation encoding capability, we propose a pretraining strategy based on degraded image reconstruction. This strategy leverages a hybrid dataset composed of paired samples with diverse degradation types. The key objective is to train the degradation extractor $\mathcal{E}$ to capture detailed degradation information from a degraded image $\boldsymbol{I}_{D}$ and compress it into a compact embedding $\boldsymbol{E}_{d}$, formulated as:
$\boldsymbol{E}_{d}=\mathcal{E}(\boldsymbol{I}_{D}).$ (6)
Subsequently, the reconstruction unit $\mathcal{R}$ takes $\boldsymbol{E}_{d}$ together with the clean content embedding $\boldsymbol{E}_{c}$ to reconstruct the degraded image $\boldsymbol{I}_{D}$. In particular, we adopt Restormer as the content extractor to derive $\boldsymbol{E}_{c}$ from the clean reference image $\boldsymbol{O}$, expressed as:
$\boldsymbol{E}_{c}=Restormer(\boldsymbol{O}),\ \ \boldsymbol{\hat{I}}_{D}=\mathcal{R}(\boldsymbol{E}_{c},\boldsymbol{E}_{d}).$ (7)
This pretraining paradigm encourages the degradation extractor to encode both the type and severity of degradation into a continuous embedding space, which then serves as a compact and informative task descriptor for downstream restoration. We further introduce a dual-component loss function $L$, which combines the $L_{1}$ norm and the Learned Perceptual Image Patch Similarity (LPIPS) metric. The combined loss function is expressed as $L(\boldsymbol{I}_{D},\boldsymbol{\hat{I}}_{D})=L_{1}+L_{LPIPS}$, striking a balance between reconstruction accuracy and perceptual quality.
III-D Degradation-aware Prompt State Space Model
To address the limitations of SSM-based image restoration—specifically the insufficient use of global context and the lack of effective degradation-aware mechanisms, we propose a novel Degradation-aware Prompt State Space Model (DP-SSM) to fully unlock the potential of SSM in All-in-One image restoration. Specifically, we propose a novel Degradation-aware Prompt Selective Scan, which incorporates degradation information into the State Space Model. This is achieved by leveraging a Degradation Embedding extracted from a pre-trained Degradation Extractor to provide supplementary degradation-related contextual information that the traditional SSM cannot capture. In addition, we modulate the parameter matrices in the state-space equations to establish a degradation-aware mechanism. This mechanism enables SSM to adapt to multi-task learning and handle task competition. To better recover local high-frequency details, we design a lightweight High-frequency Enhancement Block (HEB). This component ensures the preservation of high-frequency details and enhances the overall model performance.
Degradation-aware Prompt Selective Scan. We revisit the traditional SSM, whose parameter settings are shown in Eq.(2). Matrices $A$, $B$, $C$, and $D$ control the SSM output. The state transition matrix $A$ compresses historical information, while the input matrix $B$ maps input signals to the state space, determining their influence on historical states. The output matrix $C$ maps historical states to the observable space, reflecting their impact on the output. Matrix $D$ provides a direct pathway from input to output, similar to a residual connection. In traditional SSM, these matrices and $\Delta$ are derived from:
$A=init(),D=init(),$ (8)
$B=Linear(x),C=Linear(x),\Delta=Linear(x).$
where $init()$ is an initialization method that does not depend on the input $x$. However, such parameterization does not incorporate degradation information, which motivates us to introduce a degradation-aware prompt mechanism to guide the state-space modeling.
To integrate degradation awareness into the SSM, we first utilize a degradation extractor $\mathcal{E}$ to extract the degradation embedding $\boldsymbol{E}_{d}$ from the input image $I_{D}$. This embedding is then employed to modulate $\Delta$, the input matrix $B$, and the output matrix $C$, thereby enhancing the model’s ability to compress historical information, process input signals, and capture global context, as illustrated in Fig. 4 (a). The specific formulation is given as follows:
$\boldsymbol{E}_{d}$ $=\mathcal{E}(I_{D}),$ (9)
$\alpha_{\Delta}$ $=\mathcal{M}_{\Delta}(\boldsymbol{E}_{d}),\,\alpha_{B}=\mathcal{M}_{B}(\boldsymbol{E}_{d}),\,\alpha_{C}=\mathcal{M}_{C}(\boldsymbol{E}_{d}),$
$\Delta_{dp}$ $=\alpha_{\Delta}\cdot\Delta,\,B_{dp}=\alpha_{B}\cdot B,\,C_{dp}=\alpha_{C}\cdot C.$
Here, $\mathcal{M}_{\Delta}$, $\mathcal{M}_{B}$, and $\mathcal{M}_{C}$ are combinations of linear layers used to map the degradation embedding $\boldsymbol{E}_{d}$ into feature spaces, forming modulation vectors $\alpha_{\Delta}$, $\alpha_{B}$, and $\alpha_{C}$ for $\Delta$, $B$, and $C$, respectively. $\Delta_{dp}$, $B_{dp}$, and $C_{dp}$ are the newly modulated SSM parameters.
Finally, the formulation of the Degradation-aware State Space Model can be expressed as follows:
$\overline{A}_{dp}$ $=\exp(\Delta_{dp}A),$ (10)
$\overline{B}_{dp}$ $=(\Delta_{dp}A)^{-1}(\exp(A)-I)\cdot\Delta_{dp}B_{dp},$
$h_{i}$ $=\overline{A}_{dp_{i}}h_{i-1}+\overline{B}_{dp_{i}}x_{i},$
$y_{i}$ $=C_{dp}h_{i}+Dx_{i}.$
Through this approach, the transformation matrix $A$ and input matrix $B$ gain degradation-aware capabilities, integrating degradation cues of various types and levels during historical information compression and input signal processing, making the model more task-friendly. The output matrix $C$ controls how historical information affects the output. In traditional SSM, the $i$-th pixel can only observe the i1i-1 preceding pixels, leading to a limited global receptive field. While existing methods address this with parallel multi-directional scanning, this increases computational costs and offers limited performance gains in low-level vision tasks [32]. Instead, we retain single-directional scanning and enhance global degradation cues via $C$, achieving efficient global awareness with lower computational overhead.
High-frequency Enhancement Block.
Figure 5: We compare frequency-domain differences between clear and restored images under three configurations. The All-in-One setting shows a notable high-frequency gap compared to the Degradation-specific setting, which is significantly reduced by adding the High-frequency Enhancement Block (HEB), as shown by the blue line.
All-in-One image restoration involves recovering image components distributed across different frequency bands. While our proposed Degradation-aware Prompt State Space Model enables dynamic adaptation to degradation characteristics and partially alleviates inter-task conflicts, it still exhibits a bias toward low-frequency restoration due to the inherent global modeling nature of Mamba. As shown in Fig. 5, All-in-One models tend to underperform in reconstructing high-frequency details such as edges and textures, compared to degradation-specific models. To address this limitation, we introduce an extremely lightweight High-Frequency Enhancement Block (HEB) as a complementary module.
For input features $F$, we first extract channel-wise low-frequency features using global average pooling, then subtract them from $F$ to obtain high-frequency components, as illustrated in Fig. 4(b). Finally, the high-frequency features are added back to enhance $F$, formulated as:
${F}_{enhanced}={F}+\alpha\cdot\big({F}-\mathcal{G}({F})\big).$ (11)
Here, $\alpha$ represents a learnable scalar that controls the contribution of high-frequency enhancement, and $\mathcal{G}$ denotes the global average pooling operation. ${F}_{enhanced}$ is the enhanced feature map.
III-E Loss Function
We employ the commonly used L1 loss $\mathcal{L}_{1}$ and L2 loss $\mathcal{L}_{2}$ in low-level vision tasks. Additionally, following previous works [39, 40], we incorporate a frequency-domain loss $\mathcal{L}_{fft}$ for training. The total loss is defined as:
[…]
where $O$ and $\hat{O}$ denote the model output and the ground truth, respectively. The parameters $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are balancing factors, which we set to 1, 0.5 and 0.001, respectively.
Figure 6: Examples of Seven Types of Image Degradations in the Dataset.
TABLE I: Quantitative Comparison on Seven Datasets for All-in-One Image Restoration Task with State-of-the-Art Methods. The Best and Second-Best Results Are in Bold and Underline.
Method Lowlight Snow Rain Haze Jpeg Noise Blur Average PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM MPRNet 19.71 0.659 30.94 0.938 26.27 0.912 25.63 0.947 30.51 0.938 28.03 0.897 26.07 0.869 26.74 0.880 MIRNet 19.81 0.668 31.04 0.941 25.77 0.907 22.26 0.882 30.54 0.939 28.14 0.900 26.32 0.873 26.27 0.873 NAFNet 19.49 0.650 30.33 0.935 26.04 0.906 25.63 0.946 30.20 0.936 27.80 0.895 26.13 0.870 26.52 0.877 Restormer 19.99 0.670 31.67 0.944 26.91 0.919 26.18 0.953 30.56 0.939 28.17 0.890 26.62 0.880 27.16 0.886 MambaIR 19.81 0.664 30.43 0.935 25.52 0.901 25.90 0.944 30.38 0.937 27.91 0.895 26.07 0.869 26.57 0.878 PromptIR 19.90 0.669 31.52 0.943 26.80 0.918 26.13 0.945 30.57 0.939 28.18 0.901 26.36 0.874 27.07 0.884 IDR 19.68 0.663 31.16 0.941 26.39 0.914 26.11 0.948 30.51 0.939 28.10 0.899 26.29 0.873 26.89 0.883 NDR-Restore 19.79 0.635 30.00 0.902 26.00 0.813 27.00 0.943 30.19 0.872 27.78 0.786 25.77 0.791 26.64 0.820 OneRestore 19.77 0.657 29.95 0.931 25.37 0.900 27.74 0.962 30.10 0.934 27.65 0.890 25.90 0.866 26.64 0.877 MoCEIR 19.92 0.674 31.22 0.943 26.56 0.920 28.06 0.963 30.51 0.939 28.13 0.900 26.34 0.874 27.25 0.888 AdaIR 19.99 0.672 31.94 0.923 27.01 0.922 26.61 0.952 30.53 0.939 28.19 0.900 26.61 0.879 27.26 0.887 DPMambaIR 20.04 0.680 32.36 0.949 27.07 0.926 28.13 0.963 30.57 0.939 28.25 0.902 27.43 0.894 27.69 0.893 DPMambaIR-L 20.06 0.677 32.31 0.950 27.11 0.925 28.39 0.963 30.55 0.939 28.21 0.902 27.45 0.895 27.73 0.893
TABLE II: Quantitative Comparison on Deraining, Low-light Image Enhancement(LLIE), Deblurring and Dehazing for Degradation-Specific Image Restoration Task with State-of-the-Art Methods. The Best and Second-Best Results Are in Bold and Underline.
Deraining PSNR SSIM JORDER 26.25 0.835 PReNet 29.46 0.899 MPRNet 30.41 0.891 MAXIM 30.81 0.903 Restormer 31.46 0.904 OneRestore 29.36 0.944 MoCEIR 31.23 0.960 AdaIR 31.46 0.961 DPMambaIR 32.30 0.967
LLIE PSNR SSIM EnlightenGAN 17.61 0.653 MIRNet 24.14 0.830 URetinex-Net 19.84 0.824 MAXIM 23.43 0.863 IAT 23.38 0.809 OneRestore 22.97 0.835 MoCEIR 23.64 0.818 AdaIR 23.47 0.831 DPMambaIR 24.20 0.852
Deblurring PSNR SSIM DeepDeBlur 29.08 0.913 DeBlurGAN 28.70 0.858 DeBlurGANV2 29.55 0.934 MT-RNN 31.15 0.945 IR-SDE 30.70 0.901 OneRestore 28.76 0.915 MoCEIR 30.05 0.933 AdaIR 30.57 0.939 DPMambaIR 31.42 0.948
Dehazing PSNR SSIM GCANet 26.59 0.935 GridDehazeNet 25.86 0.944 DeHazeFormer 30.29 0.964 MAXIM 29.12 0.932 DA-CLIP 30.16 0.936 OneRestore 30.16 0.974 MoCEIR 30.62 0.975 AdaIR 30.87 0.975 DPMambaIR 30.89 0.977
IV Experiment and Analysis
We evaluate our method on both All-in-one image restoration tasks and Degradation-specific image restoration tasks. In the All-in-One Image Restoration setting, we train a single model on a dataset containing seven types of degradation. For Degradation-specific image restoration, we train separately on datasets corresponding to each type of degradation.
IV-A Experimental Settings
Training Details. Building on previous works[39, 22, 32], we develop a Mamba-based baseline for image restoration. Our method is trained using the AdamW[41] optimizer with parameters $\beta_{1}=0.9$, $\beta_{2}=0.999$, and a weight decay of $1e^{-4}$ for 300,000 iterations. The initial learning rate of $3e^{-4}$ is gradually reduced to $1e^{-6}$ through cosine annealing. Following[39], we employ a progressive training strategy, cropping image patches of size $192\times 192$. Horizontal and vertical flips are applied for data augmentation. Our experiments are conducted on a single NVIDIA RTX A100 GPU and implemented using the PyTorch platform.
Datasets. For All-in-One image restoration, we collect a large-scale mixed dataset covering seven common degradation types: haze, rain, snow, low-light, blur, noise, and JPEG compression, as shown in Fig. 6. Detailed information regarding these datasets is summarized in Table III. Additionally, degradation-specific evaluations are conducted using LOL [42] for low-light enhancement, GoPro [43] for deblurring, Rain100H [44] for deraining, and RESIDE6K [45] for dehazing.
TABLE III: Datasets for Image Restoration Tasks
Degradation Dataset Train Pairs Test Pairs
Haze RESIDE6K [45] 6000 1000
Rain Rain13K [46] 13711 -
Rain100H [44] - 100
Snow Snow100K [47] 5000 1000
Low Light LOL [42], LSRW [48] 6085 65
Blur GoPro [43] 2103 1111
Noise WaterlooED [49] 4744 -
BSD400 [50] - 400
JPEG WaterlooED [49] 4744 -
BSD400 [50] - 400
Figure 7: Qualitative comparison on Deblurring, Deraining and Dehazing for All-in-One Image Restoration Task.
Evaluation Metrics. Following previous works[39, 32], we use peak signal-to-noise ratio (PSNR) and structural similarity index measure (SSIM) to quantitatively compare different methods on the RGB channels.
Comparisons with State-of-the-art Methods. For All-in-One image restoration, we compare against eleven methods: MIRNet [51], NAFNet [52], MPRNet [53], Restormer [39], MambaIR [32], PromptIR [22], IDR [23], NDR-Restore [54], OneRestore [9], MoCEIR [55], and AdaIR [24]. The first four are general-purpose restoration models, while the latter four are designed specifically for All-in-One restoration: PromptIR uses visual prompts, OneRestore employs a pre-trained degradation encoder with cross-attention, and AdaIR leverages frequency-domain information for adaptive restoration. For degradation-specific restoration, we also include representative single-task methods such as JORDER [56] (deraining), MAXIM [57] (deblurring), EnlightenGAN [58] (low-light enhancement), DeepDeblur [59], AdaIR [24], DA-CLIP [25], among others.
IV-B Quantitative Results
All-in-one Image Restoration. We conducted two versions of our method for comparison: the normal version (DPMambaIR) and a Large version (DPMambaIR-L). The normal version adopts the degradation extractor illustrated in Fig. 3(c), while the Large version uses the architecture shown in Fig. 3(b). For subsequent evaluations, we primarily adopt the normal version unless otherwise specified. We compare our method against eleven state-of-the-art image restoration approaches. As shown in Table I, for the All-in-One image restoration task, our method achieves the best PSNR and SSIM on the aforementioned mixed datasets, outperforming existing methods. Specifically, our method achieves a PSNR of 27.69dB and an SSIM of 0.893, surpassing AdaIR by 0.43dB and 0.011, respectively. Additionally, DPMambaIR-L achieves a PSNR of 27.73dB and an SSIM of 0.893, slightly outperforming the normal version.
We achieve the best performance across all seven sub-tasks. Experimental results indicate that multi-task learning often encounters challenges in balancing performance across different tasks. For instance, AdaIR demonstrates strong performance in deraining but performs poorly in dehazing. On the other hand, OneRestore excels in dehazing but delivers suboptimal results in tasks such as deraining and deblurring. In contrast, our method effectively tackles the balance issue in multi-task learning, achieving consistently optimal performance across all sub-tasks. These findings highlight the effectiveness and superiority of our proposed method.
Degradation-Specific Image Restoration. We evaluate our method on four Degradation-specific tasks, consistently achieving superior performance across multiple datasets compared to state-of-the-art CNN-based, transformer-based, and diffusion-based approaches, as shown in Table II. On Rain100H [44], our method achieves a PSNR of 32.30dB and SSIM of 0.967, surpassing AdaIR by 0.84dB and 0.006, respectively. On the LOL dataset [42], it achieves 24.20dB PSNR and 0.852 SSIM, outperforming MIRNet by 0.06dB and 0.022. For deblurring on GoPro [43], it attains 31.42dB PSNR and 0.948 SSIM, exceeding MT-RNN by 0.27dB and 0.003. On RESIDE6K [45], it achieves 30.89dB PSNR and 0.973 SSIM, outperforming AdaIR by 0.02dB and 0.002. These results highlight the versatility and effectiveness of our method in addressing diverse degradation restoration tasks.
IV-C Qualitative Results
Fig. 7 presents the visual results of deblurring, deraining, and dehazing under the All-in-One Image Restoration task setting. Our method demonstrates superior visual performance compared to other approaches. Specifically, in deblurring, our method restores sharper edge details; in deraining, it more effectively removes rain streaks while recovering background information; and in dehazing, it focuses on regions often overlooked by other methods, resulting in more comprehensive dehazing performance. Additional visual comparisons under both the All-in-One and Degradation-specific settings are provided in the Appendix.
IV-D Ablation Study
We conduct ablation studies to evaluate our key designs: the degradation-aware prompts and High-frequency Enhancement Block (HEB) (Table IV), degradation extraction and utilization methods (Table V), and the global information supplementation strategy (Table VI).
Table IV shows the effectiveness of our core modules. The High-frequency Enhancement Block (HEB) alone improves the baseline by 0.29 dB PSNR by enhancing local details. Degradation-aware prompts ($P_{\Delta}$, $P_{B}$) bring a more significant gain of 0.64 dB, showing the benefit of adaptive modulation. Adding $P_{C}$ for global cues offers further slight improvement. The full model demonstrates the synergy between prompt-based modulation and frequency enhancement, achieving the best performance.
TABLE IV: Ablation Studies on Our Proposed Core Modules.
Method $P_{\Delta}$ $P_{B}$ $P_{C}$ HEB PSNR SSIM Baseline 26.96 0.884 (a) 27.25 0.887 (b) 27.60 0.892 (c) 27.67 0.893 (d) 27.69 0.893
Figure 8: t-SNE visualization of degradation embeddings.
Table V presents the results for degradation extraction and utilization methods. For degradation extraction, we compare three approaches: (1) “Classes”, which provides manually specified degradation types and can be regarded as a 100% accurate non-blind setting; (2) a pre-trained degradation classifier from OneRestore [9]; and (3) our fine-grained degradation representation via image reconstruction. The OneRestore extractor, optimized for classification, is insensitive to degradation strength or location, leading to poorer results. Our method surpasses the non-blind ”Classes” baseline by 0.11 dB PSNR.
Fig. 8 visualizes the t-SNE embeddings of the extracted degradation features. Significant gaps exist between different degradation types, indicating clear separability. Notably, although JPEG and noise share the same content during synthesis, the t-SNE visualization reveals that their separability is primarily driven by degradation information rather than image content.
For degradation utilization, we compare concatenation and attention-based fusion with our prompt modulation design. Our prompt-based modulation outperforms simple concatenation or attention-based fusion, showing better adaptability to local degradations.
TABLE V: Ablation Study on Different Degradation Extraction and Utilization Methods.
Comparison Projects Method PSNR SSIM Degradation Extraction Classes 27.58 0.892 OneRestore* 27.56 0.891 Ours 27.69 0.893 Degradation Insertion Concat 26.96 0.882 Attention 27.39 0.882 Ours 27.69 0.893
Table VI presents the ablation study on global information supplementation. We compare our Prompt-C with Mamba’s bi-directional scanning scheme. Compared to Mamba’s bi-directional scan, our Prompt-C not only achieves slightly better performance but is also more efficient, reducing GFLOPs by nearly 37% (Table VI). This highlights its effectiveness as a lightweight global information supplementation method.
TABLE VI: Ablation Study on Different Methods for Global Information Supplementation.
Method PSNR SSIM Params(M) GFLOPs Bi-direction Scan 27.66 0.893 31.43 244.3 Prompt Matrix C (Ours) 27.69 0.893 29.35 153.3
V Conclusion
We propose DPMambaIR, a novel All-in-One framework for image restoration capable of handling diverse degradation types. The core of our framework is a Degradation-Aware Prompt State Space Model (DP-SSM) that leverages a fine-grained degradation extractor. This design enables the dynamic integration of degradation features into the state-space modeling process, allowing for adaptive handling of complex degradation scenarios while maintaining global degradation awareness. Furthermore, a lightweight High-frequency Enhancement Block (HEB) was introduced to complement the main framework by enhancing high-frequency details restoration with negligible computational overhead. This study underscores the importance of fine-grained degradation modeling and dynamic feature modulation in advancing All-in-One image restoration frameworks. Extensive experiments on a mixed dataset containing seven degradation types show that DPMambaIR achieves the best performance, establishing it as a promising approach for unified image restoration.